24.959 Workshop in Syntax and Semantics Markéta Ceplová Syntax Paper Draft

Subject Doubling and Some Other Overt Pronouns in Czech ^{∗}

May 27, 2002

1 Introduction

Czech language has been claimed to be a Null-Subject language (Franks, 1995; Lindseth, 1996), mostly due to the pervasive availability of leaving out a pronominal subject unless they are stressed/focused. However, it has also been observed that, somewhat surprisingly for a Null-Subjectlanguage, Czech optionally allows overt subject pronouns in addition to null pro in many contexts (often without any or only slight changes in meaning). The most notable of these contexts are optional overt expletives. Apart from these, spoken Czech also has a construction where an overt pronoun optionally appears in the same clause as a full DP subject (which I will call, after Lindseth (1996), Subject Doubling construction).

In this paper I will first survey the Subject Doubling construction with the hope that once understood, it may provide us with means of looking at overt expletives and possibly also other optional overt subject pronouns in Czech. I will first describe the relevant data and distribution of the construction and then provide three possibilities of an account. I will also discuss different predictions these accounts make. I will then suggest a possible way of assimilating the optional overt expletives to this construction.

2 Subject Doubling Construction

• obligatory agreement

• separating pronoun and the full DP

• only pron + NP (never NP + NP, or NP + pron)

∗ I would like to thank Danny Fox, Sabine Iatridou, Shigeru Miyagawa, David Pesetsky, and my

classmates for many helpful comments and discussions. All remaining mistakes are my own.

3 Possible Approaches

In this section I look at XX possible structural approaches to the Subject Doublingconstruction. I sketch out a very simple basics of the structure (usually to the extent that it has been proposed) and then look at

3.1 Double Subject Structure

Lindseth (1996) in her dissertation proposes that the pronoun is base-generated separately from the related NP. The pronoun (pron) starts out in the canonical subject position, which in Lindseth’s version of split IP means [Spec,AgrP], and gets its Case checked from the Agr head with which it agrees. The NP subject stays in situ in [Spec,VP] and gets its Case licensed under goverement from the T head which heads a TP projection below AgrP and VP.

(1)

AgrP

[figure]

pron

Agr’

[figure]

Agr [figure: TP; T; VP; _{NP V}]

This proposal is not a fully specified account, though. It does not specify how the two subject arguments are related, if at all (but recall the obligatory agreement between the full DP and the pronoun) and what licenses two Nominative assigning heads^{1}. It does not provide an explanation of what is special in Czech as to provide two subject slots. As Lindseth claims, the pronoun is a true expletive. Even if true, it still does not explain why Czech should have (optional) expletives. It also doesn’t explain why this construction is limited to a personal pronoun plus an NP where the pronoun crucially precedes the NP.

A further problem surfaces with respect to the distribution of the Subject Doubling construction. If the pronoun was truly in the regular subject position, we would expect the distribution of the Subject Doubling constructions to be strictly parallel to clauses with overt subject. This, however, doesn’t seem to be entirely correct. One environment where the strict parallelism does not hold are certain embedded clauses (e.g. “when” or “if” clauses). While (2a) is perfectly acceptable (to the extent that any Subject Doubling construction is acceptable), (2b) is significantly worse.

1 One possibility would be to say that T and Agr heads have to form a chain with whatever properties

that allow this and further force the two arguments to agree

(2) a. Nemám rád, not-like.1s když when {se \atop refl} Mánička Mánička.nom všemu to-all směje. laughs

‘I don’t like it when Mánička laughs at everything’

b. ?? Nemám rád, not-like.1s když when vona she.nom {se \atop refl} Mánička Mánička.nom všemu to-all směje. laughs

‘I don’t like it when Mánička laughs at everything’

The other environment where the parallelism brakes down are sentences with topical elements (e.g. a contextually salient adverb like “yesterday”) preceding the pronoun. Both (3a) with “yesterday” preceding the subject DP, and (3b) with Subject Doubling are fine, but combining these two gives a degraded result in (3c).

(3) a. Včera yesterday {se \atop refl} Mánička Mánička.nom všemu to-all smála. laughed.fem

‘Yesterday Mánička laughed at everything’

b. Vona she.nom {se \atop refl} Mánička Mánička.nom všemu to-all smála. laughed.fem

‘Mánička laughed at everything’

c. ?? Včera yesterday vona she.nom {se \atop refl} Mánička Mánička.nom všemu to-all smála. laughed.fem

‘Mánička laughed at everything’

Given the vagueness of this proposal, it is hard to see if and how it could be improved to predict the degradation. Lindseth offers a pragmatic story based on the general pragmatic oddness or implausibility of such constructions for some expletives which share the same problem. However, that pragmatic story doesn’t seem to be applicable here since there doesn’t seem to be any strong intuitive difference in meaning between the construction with and without the overt pronoun. Whatever pragmatic approach therefore should not distinguish between them either.

3.2 Pronoun+NP as a Constituent

Another possible account for the Subject Doublingconstruction could be that the pronoun and the NP start out as a constituent and then the pronoun moves out (possibly in a way similar to clitic doubling). This structure is what Kučerová (ms.) suggests for this construction in her handout.

(4)

DP

[figure]

pron [figure: DP; _{D NP}]

Once again, this is not a well spelled-out proposal which is what would be needed to see whether or not it works, but here are some immediate advantages and problems. First of all, this structure provides a straightforward way of accounting for the agreement between the pronoun and the NP. It also provides an iteresting answer to the question

4 Conclusion

bla bla bla

References

Franks, Steven. 1995. Parameters of Slavic morphosyntax. New York – Oxford: Oxford University Press.

Kučerová, Ivona. ms. Nulový subjekt v češtině. Typologické poznámky [Null subject in Czech. Typological notes]. Presentation handout.

Lindseth, Martina. 1996. Null-Subject properties of Slavic languages: With special reference to Russian, Czech and Sorbian. Doctoral Dissertation, Indiana University.